# Unveiling m-Sharpness Through the Structure of Stochastic Gradient Noise

### v1

@@ Abstract
Sharpness‐aware minimization (SAM) has emerged as a highly effective technique for improving model generalization, but its underlying principles are not fully understood. We investigated the phenomenon known as m-sharpness, where the performance of SAM improves monotonically as the micro-batch size for computing perturbations decreases. Leveraging an extended Stochastic Differential Equation (SDE) framework, combined with an analysis of the structure of stochastic gradient noise (SGN), we precisely characterize the dynamics of various SAM variants. Our findings reveal that the stochastic noise introduced during SAM perturbations inherently induces a variance-based sharpness regularization effect. Motivated by our theoretical insights, we introduce Reweighted SAM, which employs sharpness-weighted sampling to mimic the generalization benefits of m-SAM while remaining parallelizable. Comprehensive experiments validate the effectiveness of our theoretical analysis and proposed method.
@@ 1 Introduction
We investigate the recently proposed Sharpness-Aware Minimization (SAM) (Foret et al.,, 2021), which has achieved remarkable success in various application domains (Foret et al.,, 2021; Kwon et al.,, 2021; Kaddour et al.,, 2022; Liu et al.,, 2022). It seeks flat minima by minimizing the perturbed loss $\min_{x\in\mathbb{R}^{d}}f\bigl(x+\rho\,\epsilon\bigr),$ where the perturbation $\epsilon$ is defined via
@@
m-SAM and m-sharpness. m-SAM refers to dividing a mini-batch of data into disjoint micro-batches of size $m$, and independently computing perturbations and gradients for each micro-batch, which are then combined to update the parameters. Although m-SAM was originally introduced to improve computational efficiency in distributed training, it has been observed that for $m$ below the batch size, its practical performance increases monotonically as $m$ decreases—a phenomenon known as m-sharpness (Foret et al.,, 2021; Behdin et al.,, 2022; Andriushchenko and Flammarion,, 2022). The underlying cause of this phenomenon remains an open question.
@@
Remark 1.1.
In the original formulation by Foret et al., (2021), m-SAM was defined in a multi-GPU training setup by setting $m=\text{per-GPU batch size}$, which effectively amounts to a distributed mini-batch SAM. Later, Andriushchenko and Flammarion, (2022); Behdin et al., (2022) generalized this framework to allow $m<\text{per-GPU batch size}$. Throughout this paper, in the multi-GPU setting we refer to m-SAM using the latter definition, which is inherently not parallelizable.
We extend the recent Stochastic Differential Equation (SDE) framework of Li et al., (2019); Compagnoni et al., (2023); Luo et al., (2025) by jointly tracking both $\eta$ and $\rho$ to arbitrary expansion orders, providing a unified basis for analyzing SAM and its variants. Under this framework, we derive closed-form drift terms for three unnormalized SAM (USAM) variants—n-USAM, mini-batch USAM, and m-USAM—revealing how stochastic gradient noise (SGN) drives implicit sharpness regularization and closely correlates with generalization performance. We further extend our analysis to the three normalized (vanilla) SAM variants and observe a similar noise-induced pattern, albeit without closed-form solutions. Motivated by our theory, we propose a sample‐reweighting method that uses the magnitude of the SGN as an importance measure.
@@
  • Reweighted-SAM algorithm: we introduce an adaptive reweighting mechanism that assigns larger weights to samples with higher SGN magnitudes, thereby strengthening implicit sharpness regularization; its superior generalization is confirmed through comprehensive experiments.
@@ 2 Related Works
Structure of stochastic gradient noise. In expectation, a widely accepted assumption is that the stochastic gradient serves as an unbiased estimator of the full-batch gradient. Regarding the covariance of SGN, Simsekli et al., (2019) assumed it to be isotropic. However, this view was later challenged by Xie et al., (2020) and Li et al., (2021), who argued that Simsekli et al., (2019) were actually analyzing gradient noise across different iterations rather than noise arising from mini-batch sampling. They provided extensive evidence supporting the idea that the latter can be well modeled as a multivariate Gaussian variable with an anisotropic/parameter-dependent covariance structure. Furthermore, Xie et al., (2023) conducted statistical tests on the Gaussianity to support this perspective.
@@ 3.4 SDE approximation for SAM variants
From Theorems and proposition above, we observe that mini-batch SAM (like USAM) regularizes the magnitude of the SGN. In contrast, n-SAM loses this noise-regularization effect—hence its degraded performance—whereas m-SAM amplifies it, resulting in an improved performance. One intuitive explanation is that in larger batches, these noise terms tend to cancel each other out, causing the averaged stochastic gradient to concentrate around its expectation and diminishing the contribution of SGN.
@@ 3.5 Benefits of SGN Covariance Regularization for Generalization
where $\mathrm{FIM}(x)$ denotes the empirical Fisher information matrix and $\nabla^{2}f(x)$ the Hessian matrix. Moreover, an empirical result by Xie et al., (2020) shows that for neural networks, this relationship remains approximately valid even far away from the local minima. Therefore, in the late stages of training, regularizing the trace of the SGN covariance is approximately equivalent to regularizing the trace of the Hessian, which is widely regarded as a good measure of sharpness that has been observed to correlate strongly with generalization performance (Blanc et al.,, 2020; Wen et al.,, 2022; Arora et al.,, 2022; Damian et al.,, 2022; Ahn et al.,, 2023; Tahmasebi et al.,, 2024).
@@ 4 Practical Method
In this section, we will demonstrate how to translate the theoretical insights gained from our SDE approximations into a practical method. While m-SAM [^1] In the multi-GPU setting, we use “m-SAM” to denote the scenario where $m<\text{per-GPU batch size}$, which is inherently not parallelizable as well. achieves substantially better generalization than mini-batch SAM by strengthening the regularization of the SGN magnitude, its inherently sequential nature creates a serious parallelization bottleneck (see Table 6 in Appendix J for performance and time cost). This is because it must sequentially compute the perturbation for each micro‐batch and then individually backpropagate to obtain the perturbed gradient. Building on our theoretical insights, a natural question arises:
Can we design a parallelizable algorithm that preserves the generalization advantages of m-SAM? One important observation is: if we further assume that the SGN is orthogonal to the full gradient [^2] In the squared‐norm and expectation setting, which corresponds to USAM, (16) holds automatically.—which commonly arises in signal‐to‐noise‐ratio analyses (Cao et al.,, 2022; Jelassi et al.,, 2022; Zou et al.,, 2023; Huang et al.,, 2023; Allen-Zhu and Li,, 2023; Han et al.,, 2024; Huang et al.,, 2024; Li et al., 2024a, ) —then from the following equation we observe that the norm of the stochastic gradient is directly related to the norm of the SGN:
$\|\nabla f_{i}(x)\|=\sqrt{\|\nabla f(x)\|^{2}+\|\xi_{i}\|^{2}}.$ (16)
This decomposition reveals that samples with larger gradient norms carry higher‐magnitude SGN. Drawing inspiration from importance sampling—which assigns a weight to each sample proportional to its “importance”—we adapt this idea to emphasize more important samples when computing SAM’s perturbation (2). In our framework, we quantify importance by the magnitude of the SGN: samples exhibiting larger SGN contribute more to sharpness regularization and therefore deserve greater weight in the perturbation. Specifically, we introduce an adaptive weighting mechanism in which each weight $p_{i}$ reflects the importance of sample $i$ when computing the perturbation vector. This reweighting strategy defines a probability distribution $P=\{p_{i}\}_{i\in\gamma}$ over the sampled indices $i\in\gamma$, thereby modulating the influence of each sample. We refer to the resulting algorithm as Reweighted‐SAM. The objective of Reweighted‐SAM’s perturbation is formulated as follows:
@@
Additionally, we propose using a finite‐difference method combined with Monte Carlo sampling to avoid per‐sample gradient‐norm estimation via backpropagation. The formula is as follows:
@@ 5.1 Training from scratch
We evaluate three optimization methods SGD, SAM [^3] In this section, consistent with common practice, we use “SAM” to refer to the mini-batch SAM., and Reweighted-SAM on CIFAR-10 and CIFAR-100 (Krizhevsky et al.,, 2009), training three models from scratch: ResNet-18, ResNet-50 (He et al.,, 2016), and WideResNet-28-10 (Zagoruyko and Komodakis,, 2016). We use a batch size of 128 and a cosine learning rate schedule with an initial learning rate of 0.1. SAM and Reweighted-SAM are trained for 200 epochs, while SGD is trained for 400 epochs. We apply a momentum of 0.9 and a weight decay of $5e^{-4}$, along with standard data augmentation techniques, including horizontal flipping, padding by four pixels, and random cropping.
For SAM and Reweighted-SAM, we set $\rho=0.05$ for CIFAR-10 and $\rho=0.1$ for CIFAR-100. In the case of Reweighted-SAM, we determine $\delta$ through finite difference estimation on a checkpoint before training, based on the estimated error. Specifically, we use $\delta=1e-3$ for ResNet-18 and $\delta=1e-4$ for both ResNet-50 and WideResNet-28-10. For the additional hyperparameter $\lambda$ in Reweighted-SAM, we performed a grid search over $\{0.25,0.5,1.0,2.0\}$ and found that 0.5 is a very solid default choice, achieving the best performance across all experiments.
Table 1: Test accuracy comparison on CIFAR-10 and CIFAR-100 with different optimizers. RW-SAM is the abbreviation for Reweighted-SAM.
@@
We repeated three independent experiments and reported the mean and standard deviation of the test accuracy in Tables 1 and Tables 2(a). We observe that Reweighted-SAM consistently outperforms the baselines across various models, as well as on both small and large datasets.
Analysis of Computational Overhead. Reweighted-SAM requires an additional forward pass to estimate per-sample gradient norms, leading to approximately 1/6 more training overhead compared to vanilla SAM, as a forward pass typically takes about half the time of a backward pass (Kaplan,, 2022). For a report of the additional wall‐clock time overhead, please see Table 7 in Appendix J. However, Reweighted-SAM matches the performance of m-SAM at $m=64$ without incurring its nearly two-fold training overhead (See Table 6 in Appendix J). This highlights the efficiency of Reweighted-SAM in balancing computational cost and performance.
Table 2: (a) Test accuracy on ImageNet-1k with different optimizers; (b) Test accuracy fine-tuning ViT-B/16 on CIFAR-10/100 with different optimizers. RW-SAM = Reweighted-SAM.
@@ 5.2 Fine-tuning
We fine-tune a ViT-B/16 model (Dosovitskiy et al.,, 2020), pre-trained on ImageNet-1K, on CIFAR-10 and CIFAR-100. We train for 20 epochs with an initial learning rate of 0.01. Other hyperparameters are the same as those in training from scratch. The results are summarized in Table 2(b).
@@ 5.3 Robustness to label noise
Foret et al., (2021) have shown that SAM exhibits robustness to label noise. Motivated by this, we evaluate RW-SAM maintains on CIFAR-10 with labels randomly flipped at specified noise ratios. We train a ResNet-18 and report clean test accuracies in Table 3. As the noise ratio increases, RW-SAM maintains remarkably strong performance. In particular, at an 80 % noise ratio, RW-SAM achieves a 16 % absolute accuracy improvement over SAM.
@@ 5.4 Additional experiments
We train ResNet-18 on CIFAR-100 using Reweighted-SAM to evaluate its sensitivity to different values of $\lambda$. As summarized in Table 4, Reweighted-SAM demonstrates robustness to the choice of $\lambda$ and consistently outperforms the baselines within a reasonable range.
@@
According to our theory, we compare the trace of the stochastic gradient covariance matrix at convergence for different algorithms, which, near the minimum, closely approximates the Hessian matrix (See Eq. (15)). The results, shown in Table 5, indicate that compared to SGD and SAM, Reweighted-SAM indeed converges to a minimum with a smaller stochastic gradient covariance magnitude.
@@
Table 4: Reweighted-SAM $\lambda$-sensitivity

### v2

@@ Abstract
Sharpness‐aware minimization (SAM) has emerged as a highly effective technique for improving model generalization, but its underlying principles are not fully understood. We investigated the phenomenon known as m-sharpness, where the performance of SAM improves monotonically as the micro-batch size for computing perturbations decreases. In practice, the empirical m-sharpness effect underpins the deployment of SAM in distributed training, yet a rigorous theoretical account has remained lacking. To provide a theoretical explanation for m-sharpness, we leverage an extended Stochastic Differential Equation (SDE) framework and analyze the structure of stochastic gradient noise (SGN) to characterize the dynamics of various SAM variants, including n-SAM and m-SAM. Our findings reveal that the stochastic noise introduced during SAM perturbations inherently induces a variance-based sharpness regularization effect. Motivated by our theoretical insights, we introduce Reweighted SAM (RW-SAM), which employs sharpness-weighted sampling to mimic the generalization benefits of m-SAM while remaining parallelizable. Comprehensive experiments validate the effectiveness of our theoretical analysis and proposed method.
@@ 1 Introduction
We investigate the recently proposed Sharpness-Aware Minimization (SAM) (Foret et al.,, 2021), which has achieved remarkable success in various application domains (Foret et al.,, 2021; Kwon et al.,, 2021; Kaddour et al.,, 2022; Liu et al., 2022a, ). It seeks flat minima by minimizing the perturbed loss $\min_{x\in\mathbb{R}^{d}}f\bigl(x+\rho\,\epsilon\bigr),$ where the perturbation $\epsilon$ is defined via
@@
m-SAM and m-sharpness. m-SAM refers to dividing a mini-batch of data into disjoint micro-batches of size $m$, and independently computing perturbations and gradients for each micro-batch, which are then combined to update the parameters. It has been widely observed that the practical performance of m-SAM improves monotonically as $m$ decreases, a phenomenon known as m-sharpness (Foret et al.,, 2021; Behdin et al.,, 2022; Andriushchenko and Flammarion,, 2022). It is worth noting that when $m$ is smaller than the batch size, the perturbation must be computed sequentially across micro-batches, which cannot be parallelized and therefore introduces substantial additional computational overhead, although smaller $m$ typically leads to improved generalization performance.
@@
In synchronous data-parallel (multi-GPU) training with $D$ devices and per-device batch size $b$, SAM is typically implemented by computing the perturbation locally on each device using its own data and then aggregating the perturbed gradients across devices; this implementation is exactly an instance of m-SAM with $m=b$. The local-perturbation design avoids an additional cross-device synchronization in the inner perturbation step (only the outer gradient aggregation requires all-reduce), thereby reducing per-step communication overhead. The empirical observation of m-sharpness has thus become a practical cornerstone for deploying SAM at scale.
To provide a theoretical explanation for m-sharpness, we extend the recent Stochastic Differential Equation (SDE) framework of Li et al., (2019); Compagnoni et al., (2023); Luo et al., (2025) by jointly tracking both $\eta$ and $\rho$ to arbitrary expansion orders, providing a unified basis for analyzing SAM and its variants. Under this framework, we derive closed-form drift terms for three unnormalized SAM (USAM) variants—n-USAM, mini-batch USAM, and m-USAM—revealing how stochastic gradient noise (SGN) drives implicit sharpness regularization and closely correlates with generalization performance. We further extend our analysis to the three normalized (vanilla) SAM variants and observe a similar noise-induced pattern, albeit without closed-form solutions. Motivated by our theory, we propose a sample‐reweighting method that uses the magnitude of the SGN as an importance measure.
@@
  • Reweighted SAM algorithm: we introduce an adaptive reweighting mechanism that assigns larger weights to samples with higher SGN magnitudes, thereby strengthening implicit sharpness regularization; its superior generalization is confirmed through comprehensive experiments.
@@ 2 Related Works
Sharpness-Aware Minimization. Sharpness-Aware Minimization (Foret et al.,, 2021) has attracted increasing attention due to its consistent improvements in generalization across a wide range of tasks. A growing body of work has been devoted to analyzing and enhancing SAM, including investigations into its generalization principles (Andriushchenko and Flammarion,, 2022; Möllenhoff and Khan,, 2022; Wen et al.,, 2022, 2023; Agarwala and Dauphin,, 2023; Springer et al.,, 2024; Luo et al.,, 2025) and convergence properties (Khanh et al.,, 2024; Oikonomou and Loizou,, 2025), exploring its applications in various domains, and developing algorithmic variants to further improve both generalization (Kwon et al.,, 2021; Zhuang et al.,, 2022; Liu et al., 2022b, ; Kim et al.,, 2022; Li et al., 2024c, ; Wu et al.,, 2024; Tahmasebi et al.,, 2024; Li et al., 2024b, ; Li et al.,, 2025) and computational efficiency (Du et al.,, 2021; Liu et al., 2022a, ; Du et al.,, 2022; Mordido et al.,, 2023; Tan et al.,, 2024; Xie et al.,, 2024).
@@
Structure of stochastic gradient noise. In expectation, a widely accepted assumption is that the stochastic gradient serves as an unbiased estimator of the full-batch gradient (Jastrzebski et al.,, 2017; Zhu et al.,, 2018; HaoChen et al.,, 2021; Ziyin et al.,, 2021). Regarding the covariance of SGN, Simsekli et al., (2019) assumed it to be isotropic. However, this view was later challenged by Xie et al., (2020) and Li et al., (2021), who argued that Simsekli et al., (2019) were actually analyzing gradient noise across different iterations rather than noise arising from mini-batch sampling. They provided extensive evidence supporting the idea that the latter can be well modeled as a multivariate Gaussian variable with an anisotropic/parameter-dependent covariance structure. Furthermore, Xie et al., (2023) conducted statistical tests on the Gaussianity to support this perspective.
@@ 3.3 SDE approximation for USAM variants
Under Assumption 3.1 and mild regularity conditions, the solution of the following SDE (10) is an order-$(1,1)$ weak approximation of the discrete update of n-USAM (7) with batch size $|\gamma|$:
@@ 3.4 SDE approximation for SAM variants
From Theorems and proposition above, we observe that mini-batch SAM (like USAM) regularizes the magnitude of the SGN. In contrast, n-SAM loses this noise-regularization effect, leading to degraded performance, whereas m-SAM amplifies it and thus achieves improved results. One intuitive explanation is that in larger batches, these noise terms tend to cancel each other out, causing the averaged stochastic gradient to concentrate around its expectation and diminishing the contribution of SGN.
@@ 3.5 Benefits of SGN Covariance Regularization for Generalization
where $\mathrm{FIM}(x)$ denotes the empirical Fisher information matrix and $\nabla^{2}f(x)$ the Hessian matrix. Moreover, an empirical result by Xie et al., (2020) shows that for neural networks, this relationship remains approximately valid even far away from the local minima. Therefore, in the late stages of training, regularizing the trace of the SGN covariance is approximately equivalent to regularizing the trace of the Hessian, which is widely regarded as a good measure of sharpness that has been observed to correlate strongly with generalization performance (Keskar et al.,, 2017; Blanc et al.,, 2020; Wen et al.,, 2022; Arora et al.,, 2022; Damian et al.,, 2022; Ahn et al.,, 2023; Tahmasebi et al.,, 2024).
@@ 4 Practical Method
In this section, we will demonstrate how to translate the theoretical insights gained from our SDE approximations into a practical method. While m-SAM [^1] In the multi-GPU setting, we use “m-SAM” to denote the scenario where $m<\text{per-device batch size}$, which is inherently not parallelizable as well. achieves substantially better generalization than mini-batch SAM by strengthening the regularization of the SGN magnitude, its inherently sequential nature creates a serious parallelization bottleneck (see Table 7 in Appendix J for performance and time cost). This is because it must sequentially compute the perturbation for each micro‐batch and then individually backpropagate to obtain the perturbed gradient. Building on our theoretical insights, a natural question arises:
Can we design a parallelizable algorithm that preserves the generalization advantages of m-SAM? One important observation is: if we further assume that the SGN is approximately orthogonal to the full gradient, or that the norm of the full gradient is negligible compared to that of the SGN (as often occurs in the late stages of training when the model approaches convergence), which commonly arises in signal-to-noise-ratio analyses (Cao et al.,, 2022; Jelassi et al.,, 2022; Zou et al.,, 2023; Huang et al.,, 2023; Allen-Zhu and Li,, 2023; Han et al.,, 2024; Huang et al.,, 2024; Li et al., 2024a, ), then from the following equation we observe that the norm of the stochastic gradient is directly related to the norm of the SGN:
$\|\nabla f_{i}(x)\|\approx\sqrt{\|\nabla f(x)\|^{2}+\|\xi_{i}\|^{2}}.$ (16)
This decomposition reveals that samples with larger gradient norms carry higher‐magnitude SGN. Drawing inspiration from importance sampling, which assigns a weight to each sample proportional to its “importance”, we adapt this idea to emphasize more important samples when computing SAM’s perturbation (2). In our framework, we quantify importance by the magnitude of the SGN: samples exhibiting larger SGN contribute more to sharpness regularization and therefore deserve greater weight in the perturbation. Specifically, we introduce an adaptive weighting mechanism in which each weight $p_{i}$ reflects the importance of sample $i$ when computing the perturbation vector. This reweighting strategy defines a probability distribution $P=\{p_{i}\}_{i\in\gamma}$ over the sampled indices $i\in\gamma$, thereby modulating the influence of each sample. We refer to the resulting algorithm as Reweighted SAM (RW-SAM). The objective of RW-SAM’s perturbation is formulated as follows:
@@
Additionally, we propose using a finite‐difference method combined with Monte Carlo sampling to avoid per‐sample gradient‐norm estimation via backpropagation. The formula is as follows (see derivation in Section G):
@@ 5.1 Training from scratch
We evaluate three optimization methods SGD, SAM [^2] In this section, consistent with common practice, we use “SAM” to refer to the mini-batch SAM., and RW-SAM on CIFAR-10 and CIFAR-100 (Krizhevsky et al.,, 2009), training three models from scratch: ResNet-18, ResNet-50 (He et al.,, 2016), and WideResNet-28-10 (Zagoruyko and Komodakis,, 2016). We use a batch size of 128 and a cosine learning rate schedule with an initial learning rate of 0.1. SAM and RW-SAM are trained for 200 epochs, while SGD is trained for 400 epochs. We apply a momentum of 0.9 and a weight decay of $5e^{-4}$, along with standard data augmentation techniques, including horizontal flipping, padding by four pixels, and random cropping.
For SAM and RW-SAM, we set $\rho=0.05$ for CIFAR-10 and $\rho=0.1$ for CIFAR-100. In the case of RW-SAM, we determine $\delta$ through finite difference estimation on the model before training, based on the estimated error. Specifically, we use $\delta=1e-3$ for ResNet-18 and $\delta=1e-4$ for both ResNet-50 and WideResNet-28-10. For the additional hyperparameter $\lambda$ in RW-SAM, we performed a grid search over $\{0.25,0.5,1.0,2.0\}$ on a validation set and found that 0.5 consistently yielded strong performance across experiments.
Table 1: Test accuracy comparison on CIFAR-10 and CIFAR-100 with different optimizers.
@@
We repeated three independent experiments and reported the mean and standard deviation of the test accuracy in Table 1 and Table 2(a). We observe that RW-SAM consistently outperforms the baselines across various models, as well as on both small and large datasets.
Analysis of Computational Overhead. RW-SAM requires an additional forward pass to estimate per-sample gradient norms, leading to approximately 1/6 more training overhead compared to vanilla SAM, as a forward pass typically takes about half the time of a backward pass (Kaplan,, 2022). For a report of the additional wall‐clock time overhead, please see Table 8 in Appendix J. However, RW-SAM matches the performance of m-SAM at $m=64$ without incurring its nearly two-fold training overhead (See Table 7 in Appendix J). This highlights the efficiency of RW-SAM in balancing computational cost and performance.
Table 2: (a) Test accuracy on ImageNet-1k with different optimizers; (b) Test accuracy fine-tuning ViT-B/16 on CIFAR-10/100 with different optimizers.
@@ 5.2 Fine-tuning
We fine-tune a ViT-B/16 model (Dosovitskiy et al.,, 2020), pre-trained on ImageNet-1K, on CIFAR-10 and CIFAR-100. We train for 20 epochs with an initial learning rate of 0.01. Other hyperparameters are the same as those in training from scratch. The results are summarized in Table 2(b). We also fine-tune a pretrained DistilBERT model (Sanh et al.,, 2019) on the GLUE benchmark (Wang et al.,, 2018). We use AdamW (Loshchilov and Hutter,, 2019) as the base optimizer. The detailed hyperparameter settings are provided in Table 10 of Appendix J, and the results are presented in Table 3. We observe that RW-SAM consistently outperforms SAM in both fine-tuning experiments.
Table 3: Performance comparison on GLUE tasks using different optimizers.
Optimizer CoLA MNLI MRPC QNLI QQP RTE SST-2 STS-B Average
AdamW 0.538 0.825 0.900 0.884 0.868 0.628 0.914 0.864 0.804
SAM 0.517 0.824 0.900 0.894 0.871 0.610 0.911 0.870 0.800
RW-SAM 0.560 0.826 0.904 0.896 0.872 0.625 0.915 0.870 0.809
@@ 5.3 Robustness to label noise
Foret et al., (2021) have shown that SAM exhibits robustness to label noise. Motivated by this, we evaluate RW-SAM’s performance on CIFAR-10 with labels randomly flipped at specified noise ratios. We train a ResNet-18 and report clean test accuracies in Table 4. As the noise ratio increases, RW-SAM maintains remarkably strong performance. In particular, at an 80 % noise ratio, RW-SAM achieves a 16 % absolute accuracy improvement over SAM.
@@ 5.4 Additional experiments
We train ResNet-18 on CIFAR-100 using RW-SAM to evaluate its sensitivity to different values of $\lambda$. As summarized in Table 5, RW-SAM demonstrates robustness to the choice of $\lambda$ and consistently outperforms the baselines within a reasonable range.
@@
According to our theory, we compare the trace of the stochastic gradient covariance matrix at convergence for different algorithms, which, near the minimum, closely approximates the Hessian matrix (See Eq. (15)). The results, shown in Table 6, indicate that compared to SGD and SAM, RW-SAM indeed converges to a minimum with a smaller stochastic gradient covariance magnitude.
@@
Table 5: RW-SAM $\lambda$-sensitivity